OpenStudy (anonymous):

find the equation of the tangent line to the graph f(x)=3x^3-10

OpenStudy (phi):

any particular point on the curve?

OpenStudy (anonymous):

what is the slope of f(x)? Slope is the first derivative, right?

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

differentiate f(x) = 3x^3 - 10... f ' (x) = 3(3)x^2 = 9x^2

OpenStudy (anonymous):

Ok. I forgot to write "at the point (2,14)" how does that change it?

OpenStudy (anonymous):

YES! You need to know where along f(x) you want to find that tangent line. Any tangent to that f(x) curve will have the slope given by f ' (x), but in order to actually write a line equation, typically you write it for one specific line equation, in other words, tangent at some specific point along f(x) defined by the x-value.

OpenStudy (anonymous):

So, at that point (2,14), what is f ' (x) = f ' (2) = ?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

is that the slope?

OpenStudy (anonymous):

No, you need to evaluate that first derivative using x = 2 f ' (x) = 9x^2 f ' (2) = 9(2)^2 = 9(4) = 36

OpenStudy (anonymous):

That is the slope of the tangent line to f(x) at the point along f(x) where x = 2

OpenStudy (anonymous):

ok. so the eqation is f'(x)=36x?

OpenStudy (anonymous):

So now you have a slope of 36 and a point (2,14). You can write a line equation using the point-slope form. y - y1 = m(x - x1) using point (2,14) as (x1, y1) and using slope m = f ' (x) = f '(2) = 36

OpenStudy (anonymous):

oh! ok! I get it. thank you

OpenStudy (anonymous):

no, f ' (x) is the derivative of f(x).... in general, it is just the expression f ' (x) = 9x^2 You can evaluate f ' (x) for any x you want... here, you need to do it at x = 2

OpenStudy (anonymous):

For curves, the slope is the first derivative which is a function, not a single number. Lines have constant slope, so they have a slope m. Curves have a slope that changes with different x values, so it's an expression using x. But you can solve for slope at any SPECIFIC x by plugging in that x value into the first derivative expression.

OpenStudy (anonymous):

ok, i got y=36x-58...

OpenStudy (anonymous):

y=36x-58 is correct.

OpenStudy (anonymous):

you can check it... is (2,14) a solution to y = 36x - 58? 14 = 36(2) - 58 yes, therefore (2,14) is on the tangent line Is (2,14) on the curve f(x)? It should be... given in the problem. Is the slope of the tangent line at x = 2 equal to the derivative of f(x), f ' (x), at f'(x) = 2? Slope of tangent line is 36 everywhere. Slope of f(x) is f ' (x) = 9x^2, and at f ' (2), it is 36. So yes :) It all works out! Yay!!

OpenStudy (anonymous):

Awesome! Thanks soooooooo much!

OpenStudy (anonymous):

Glad to help :) Understand the idea? 1) Find the slope expression using first derivative. 2) Evaluate first derivative at the specific x value to get the slope value (a number). 3) Use point and slope to write equation of tangent line at that particular point. 4) Check work :)

OpenStudy (anonymous):

yup. i got it. thanks :D

OpenStudy (anonymous):

Great! Works for any question with "find equation of tangent line to some f(x) at point (x1, y1)" Good luck!

OpenStudy (anonymous):

one more quetion. The derivative of theta is zero, right?

OpenStudy (anonymous):

uh :) I don't know... is there more to the problem? Theta is often used as a variable for an angle. All alone, it doesn't mean anything specific.

OpenStudy (anonymous):

No, i was working on a second problem. it was the same question but for f(theta)=4Sin(theta)-(theta).

OpenStudy (anonymous):

oh wait. that would make the answer 1, right?

OpenStudy (anonymous):

just for the theta part.

OpenStudy (anonymous):

you have to take the derivative of the whole right side... but yes, it looks like it would be "1" f(theta) = theta is just like f(x) = x, so f '(x) = 1 Just don't forget the 4sin(theta) part ;)

OpenStudy (anonymous):

of course :) thanks